OpenStudy (anonymous):

f(x)=(x−4)/(5x+2). Find an equation for the tangent line to the graph of f at x=1. Tangent line: y =

hartnn (hartnn):

do you know how to find slope of tangent at x=1 ?

OpenStudy (anonymous):

ya its 22/49

OpenStudy (anonymous):

i should say thats the slope of the normal line

hartnn (hartnn):

hmm..i assume thats correct. so you just need 1 point now to get equation of line. when x=1, y=.. ? (put x=1, here y=f(x)=(x−4)/(5x+2).)

hartnn (hartnn):

the point will be (1,f(1)) = (1,(1−4)/(5(1)+2))

OpenStudy (anonymous):

u got -3/7

hartnn (hartnn):

you have slope = 22/49 point = (1,-3/7) can u find equation of line ?

OpenStudy (anonymous):

y=(22/49)x-(22/49)+(3/7)

hartnn (hartnn):

y=(22/49)x-(22/49)-(3/7)

OpenStudy (anonymous):

ya thts what i got

OpenStudy (anonymous):

wait why -3/7

OpenStudy (anonymous):

ah ok i see what i did my bad

hartnn (hartnn):

good :)

OpenStudy (anonymous):

messed it up in my head sorry and thank you

hartnn (hartnn):

welcome ^_^